# Myhill-Nerode for hypergraphs and an application to gain-graphic matroids

### v2

@@ 4. Amalgams
sec:amalgams
Let $M_{1}$ and $M_{2}$ be matroids with ground sets $E_{1}$ and $E_{2}$, rank functions $r_{1}$ and $r_{2}$, and closure operators $\operatorname{cl}_{1}$ and $\operatorname{cl}_{2}$. Let $\ell$ be $E_{1}\cap E_{2}$. We assume that $M_{1}|\ell=M_{2}|\ell$ and we denote this shared restriction by $N$. If $M$ is a matroid on the ground set $E_{1}\cup E_{2}$ and $M|E_{i}=M_{i}$ for $i=1,2$, then we say $M$ is an amalgam of $M_{1}$ and $M_{2}$.
A matroid is modular if $r(F)+r(F^{\prime})=r(F\cap F^{\prime})+r(F\cup F^{\prime})$ whenever $F$ and $F^{\prime}$ are flats. Assume that $N$ is a modular matroid. Then [Oxl11]*Theorem 11.4.10 tells us that we can obtain an amalgam of $M_{1}$ and $M_{2}$ by setting the rank of any subset $X\subseteq E_{1}\cup E_{2}$ to be
[…]
The resulting matroid is the proper amalgam of $M_{1}$ and $M_{2}$. We refer to the set $\ell$ as the amalgam base and we denote the proper amalgam by $\mathrm{Amal}_{\ell}(M_{1},M_{2})$, or $\mathrm{Amal}(M_{1},M_{2})$ if the amalgam base is clear from the context. We can easily check that every rank-$2$ matroid is modular. The following result is slightly more powerful than Proposition 4.1 in [MNW18] because it allows the matroids to be non-simple.
@@
We use $N$ to denote $M_{1}|\ell=M_{2}|\ell$. Assume $X\cap E_{1}$ is dependent in $M_{1}$. Then $X\cap E_{1}$ is dependent in $\mathrm{Amal}(M_{1},M_{2})$ since $\mathrm{Amal}(M_{1},M_{2})|E_{1}=M_{1}$. Thus $X$ is dependent in $\mathrm{Amal}(M_{1},M_{2})$. By symmetry we conclude that if $X\cap E_{1}$ is dependent in $M_{1}$ or if $X\cap E_{2}$ is dependent in $M_{2}$, then $X$ is dependent in $\mathrm{Amal}(M_{1},M_{2})$. Henceforth we assume $X\cap E_{1}$ is independent in $M_{1}$ and $X\cap E_{2}$ is independent in $M_{2}$.
@@
Let $Y$ be $X\cup y$. Because $y$ is in $\operatorname{cl}_{1}(X-E_{2})$ it follows that $r_{1}(X\cap E_{1})=r_{1}(Y\cap E_{1})$. Similarly, $r_{2}(X\cap E_{2})=r_{2}(Y\cap E_{2})$. If $X\cap\ell=\emptyset$ then $0=|X\cap\ell|=r_{N}(Y\cap\ell)-1$. Now assume that $X\cap\ell=\{x\}$. Since $\{x,y\}$ is independent in $N$, we have $r_{N}(Y\cap\ell)=2$, so $|X\cap\ell|=r_{N}(Y\cap\ell)-1$ holds in either case. Now we see that
@@
Assume $y$ and $y^{\prime}$ are distinct elements of $Y\cap\ell$. If $\{y,y^{\prime}\}$ is dependent then we can assume without loss of generality that $y$ is not in $X$, or else $X\cap\ell$ is dependent in $N$. Since $Y$ contains no loops, it follows that $\{y,y^{\prime}\}$ is a circuit of $N$. This means that $y$ is in $\operatorname{cl}_{N}((Y-y)\cap\ell)$, and we have contradiction to Claim 4.1.1. Therefore $\{y,y^{\prime}\}$ is independent.
We know that $Y-X\subseteq\ell$ is non-empty so $Y\cap\ell$ contains at least one element. Let $y$ be such an element. Assume $Y\cap\ell$ contains three distinct elements and let $y_{1}$ and $y_{2}$ be distinct elements in $(Y\cap\ell)-y$. Then $\{y_{1},y_{2}\}$ is independent by the previous paragraph, so $\{y_{1},y_{2}\}$ spans $N$. This means that $y$ is in $\operatorname{cl}_{N}((Y-y)\cap\ell)$, a contradiction. We conclude that $|Y\cap\ell|<3$. ∎
Assume $X\cap\ell$ contains distinct elements $x$ and $x^{\prime}$. Then $\{x,x^{\prime}\}$ is independent by Claim 4.1.2. In this case $\{x,x^{\prime}\}$ spans $N$. We can let $y$ be an element in YXY-X, and now $y$ is in $\operatorname{cl}_{N}((Y-y)\cap\ell)$, contradicting Claim 4.1.1. Hence $|X\cap\ell|$ is in $\{0,1\}$.
@@
Assume that $X\cap\ell$ is non-empty, and therefore contains a single element $x$. Without loss of generality, we assume that $x=y^{\prime}$. Now $X\cap E_{1}=(Y-y)\cap E_{1}$, so $\operatorname{cl}_{1}(X\cap E_{1})$ contains both $x$ and $y$ by Claim 4.1.1. As $\{x,y\}$ is independent in $N$ by Claim 4.1.2, it follows that $\operatorname{cl}_{1}(X\cap E_{1})$ contains $\ell$, contradicting our earlier assumption. Therefore $X\cap\ell$ is empty and $Y\cap E_{1}=(X\cap E_{1})\cup\{y,y^{\prime}\}$.
@@
lem:amalgam-is-biased
For $i=1,2$, let $\Omega_{i}=(G_{i},\mathcal{B}_{i})$ be a biased graph without balanced loops and assume that $V(G_{1})\cap V(G_{2})=\{u,v\}$. Assume also that in $\Omega_{1}$ and $\Omega_{2}$, every vertex is incident with an unbalanced loop, and that each vertex $w$ has two distinct neighbours $w_{1}$ and $w_{2}$ such that for each $i=1,2$, there is an unbalanced $2$-edge cycle containing $w$ and $w_{i}$. Let $\Omega=(G,\mathcal{B})$ be a biased graph such that $G$ has no isolated vertices and $\mathrm{Amal}(F(\Omega_{1}),F(\Omega_{2}))$ is equal to $F(\Omega)$. Then $G$ is isomorphic to $G_{1}\oplus_{uv}G_{2}$.
@@
Let $L$ be a set containing exactly one unbalanced loop incident with each vertex in $G_{1}\oplus_{uv}G_{2}$. Then $L\cap E_{i}$ is a basis of $F(\Omega_{i})$ for each $i=1,2$. It is easy to verify, using Proposition 4.1, that $L$ is independent in $\mathrm{Amal}(F(\Omega_{1}),F(\Omega_{2}))$. Let $e$ be an arbitrary element of $\mathrm{Amal}(F(\Omega_{1}),F(\Omega_{2}))$ that is not in $L$. We can assume without loss of generality that $e$ is an edge of $\Omega_{1}$. Either $e$ is an unbalanced loop that has a common vertex with a loop in $L$, or it is non-loop edge of $\Omega_{1}$. In these cases, $e$ is in a $2$- or $3$-element circuit of $F(\Omega_{1})$ that is contained in $L\cup e$. This circuit is also a circuit of $\mathrm{Amal}(F(\Omega_{1}),F(\Omega_{2}))$. It now follows that $L$ spans $\mathrm{Amal}(F(\Omega_{1}),F(\Omega_{2}))$, and is therefore a basis of this matroid. Let $x$ and $y$ be an arbitrary pair of distinct vertices contained in an unbalanced $2$-edge cycle of either $\Omega_{1}$ or $\Omega_{2}$. The edges from the $2$-edge cycle along with unbalanced loops incident with $x$ and $y$ form a long line of either $F(\Omega_{1})$ or $F(\Omega_{2})$, and it is not difficult to see that this line is also a line in $\mathrm{Amal}(F(\Omega_{1}),F(\Omega_{2}))$. Now the hypotheses imply that each element of $L$ is in two distinct long lines of $\mathrm{Amal}(F(\Omega_{1}),F(\Omega_{2}))$. Now Proposition 2.1 implies that the elements of $L$ are all unbalanced loops in $\Omega$. Since $L$ is independent in $F(\Omega)$, we have no more than one element of $L$ incident with any given vertex of $G$, for any two loops incident with the same vertex form a handcuff. If there exists some vertex of $G$ not incident with a loop in $L$, then this vertex is incident with an edge $e$ (since $G$ has no isolated vertices), and now $e$ is not spanned by $L$. This is a contradiction, so every vertex of $G$ is incident with exactly one loop in $L$. This induces a bijection between the vertices of $G_{1}\oplus_{uv}G_{2}$ and $G$. Any edge of $G_{1}\oplus_{uv}G_{2}$ that is not in $L$ is incident with at most two elements of $L$, and this shows that the bijection between the vertices of $G_{1}\oplus_{uv}G_{2}$ and $G$ gives us an isomorphism between the two graphs. ∎
@@ 5. Uniformly locally finite groups
Figure 2. The graph $H_{n,N}\setminus H_{n,N}^{*}$. fig:HnN
@@
all in $\Gamma$ satisfying condition $(\dagger)$ in Definition 5.2. Let $\sigma_{N}$ be $\sigma(\Gamma,\{a_{1},\ldots,a_{n}\},s_{N},\mathfrak{M}_{N},\{d_{i,j}\},\{q_{i}\})$, the $\Gamma$-gaining of $H_{n,N}$. Then, for ease of notation, for each $N\geq 1$, let $H_{N}=(H_{n,N},\sigma_{N})$, and let $\Omega_{N}$ be the associated biased graph. Note that any $H_{N}$ contains $G_{\ell,N}$, and the $\{G_{\ell,N}\}$ are all isomorphic as $\Gamma$-gained graphs. So, we could arrange that the $H_{n}$ are all sub-graphs of some larger graph $\mathbb{G}$, and that the $G_{\ell,N}$ are actually equal, and for any $N,N^{\prime}$ distinct, $H_{N}\cap H_{N^{\prime}}=G_{\ell,N}$. Thus, given any two $H_{N},H_{N^{\prime}}$, we can construct both the matroid amalgam $\mathrm{Amal}(F(H_{N}),F(H_{N^{\prime}}))$ over the base $F(\ell_{N})=F(\ell_{N^{\prime}})$, and the graph amalgam $H_{n,N}\oplus_{\delta_{1},\delta_{2}}H_{n,N^{\prime}}$ over the base $\{\delta_{1},\delta_{2}\}=\{\delta_{1}^{\prime},\delta_{2}^{\prime}\}$. Throughout the proof, when we amalgamate it will always be over these bases. By Lemma 5.5, if the class of $\Gamma$-gain-graphic matroids is $\operatorname{CMSO}_{1}$-definable, there is a finite partition of $\omega$ such that for any $N,N^{\prime}$, whether $\mathrm{Amal}(F(H_{N}),F(H_{N^{\prime}}))$ is a $\Gamma$-gain-graphic matroid depends only on which classes in the partition $N$ and $N^{\prime}$ belong to. Thus, for a contradiction it suffices to show that the amalgam $\mathrm{Amal}(F(H_{N}),F(H_{N^{\prime}}))$ is a $\Gamma$-gain matroid if and only if $N=N^{\prime}$ for all $N,N^{\prime}\geq 1$. There are two directions to this.
@@
and conclude $\tau(B_{i,j})=\tau(A_{j})=h_{j}$. Now, let $\eta:\{1,\ldots,N\}\rightarrow\{h_{1},\ldots,h_{n}\}\cup\{\mathrm{Id}\}$ be a string of length at most $N$ in the $h_{i}$. There is a path $P_{\eta}$ from $\gamma_{1}$ to $\gamma_{N+1}$ such that for each $k$, the edge $P_{\eta}$ takes between $\gamma_{k}$ and $\gamma_{k+1}$ is mapped to $\eta(k)$ by $\tau$. Let $C_{\eta}$ be the cycle obtained from $P_{\eta}$ by appending $\gamma_{N+1}\beta_{2}\alpha_{2}\ A_{s}^{-1}\ \alpha_{1}\ K_{1}^{-1}\ \beta_{1}\ K_{2}\ \gamma_{1}$. Thus, $\tau(C_{\eta})=\tau(P_{\eta})h_{s}^{-1}$. Since $f_{\{a_{i}\}}(s)=N$, there exists $\eta$ a string of length $N$ such that $\sigma_{N}(C_{\eta})=\mathrm{Id}$, but not one of smaller length. So there exists $\eta$ of length $N$ such that $\tau(C_{\eta})=\mathrm{Id}$ but not one of smaller length. In other words, $f_{\{h_{i}\}}(h_{s})=N$. Symmetrically, if we let $\tau^{\prime}(C_{j})=h_{j}^{\prime}$ for all $j$, and $\tau^{\prime}(C_{s})=h_{s}^{\prime}$, then $f_{\{h_{i}^{\prime}\}}(h_{s}^{\prime})=N^{\prime}$. But the edges in $\mathcal{C}$ are all in the amalgamation base, so $\tau$ and $\tau^{\prime}$ agree on $\mathcal{C}$, so
@@
Now that we have this identity, we want to show $F(H_{n,N}\oplus_{\ell_{N}}H_{n,N^{\prime}},\sigma_{N}\cup\tau^{\prime})$ is the matroid $\mathrm{Amal}(F(H_{N}),F(H_{N^{\prime}}))=\mathrm{Amal}(F(H_{N}),F(H_{n,N^{\prime}},\tau^{\prime}))$, by Lemma 4.4. We must prove that this data satisfies its conditions. Note (i), (ii) are immediate from the definition of $\tau^{\prime}$, and (iii) is satisfied by $Q_{N+6}$ and $Q_{N+7}$, so there is only one condition of Lemma 4.4 remaining: to show that for any paths $P\subseteq H_{n,N}$, $P^{\prime}\subseteq H_{n,N^{\prime}}$, both starting at $\delta_{1}$ and ending at $\delta_{2}$, and $\sigma_{N}(P)=\tau^{\prime}(P^{\prime})$, there is some edge $E$ from $\delta_{1}$ to $\delta_{2}$ such that $\sigma_{N}(E)=\sigma_{N}(P)=\tau^{\prime}(P^{\prime})$. Fix such $P,P^{\prime}$. Their union is a balanced cycle, so they must lie in $H^{*}_{n,N}$ and $H^{*}_{n,N^{\prime}}$, respectively. Then, in particular, $f_{\{a_{i}\}}(P)=f_{\{a_{i}\}}(P^{\prime})$. By inspection $f_{\{a_{i}\}}(P)\leq 3N+1$. If $P^{\prime}$ passes through $K_{1}^{\prime}$ or $K_{2}^{\prime}$, $f_{\{a_{i}\}}(P^{\prime})\geq 3N+2$. Therefore, $P^{\prime}$ passes through neither $K_{1}^{\prime}$ nor $K_{2}^{\prime}$. Hence, $\tau^{\prime}(P^{\prime})\in\{\mathrm{Id},a_{1},\ldots,a_{n},s\}$, and so there is some edge $E\in\mathcal{A}$ such that $\sigma_{N}(E)=\sigma_{N}(P)=\tau^{\prime}(P^{\prime})$. The conditions of the Lemma are satisfied, so $\mathrm{Amal}(F(H_{N}),F(H_{N^{\prime}}))=F(H_{n,N}\oplus_{\ell_{N}}H_{n,N^{\prime}},\sigma_{N}\cup\tau^{\prime})$ is a $\Gamma$-gain graphic matroid. ∎

### v3

@@ 4. Amalgams
Let $M_{1}$ and $M_{2}$ be matroids with ground sets $E_{1}$ and $E_{2}$, rank functions $r_{1}$ and $r_{2}$, and closure operators $\operatorname{cl}_{1}$ and $\operatorname{cl}_{2}$. Let $\ell$ be $E_{1}\cap E_{2}$. We assume that $M_{1}|\ell=M_{2}|\ell$ and we denote this shared restriction by NN. If $M$ is a matroid on the ground set $E_{1}\cup E_{2}$ and $M|E_{i}=M_{i}$ for $i=1,2$, then we say $M$ is an amalgam of $M_{1}$ and $M_{2}$.
A matroid is modular if $r(F)+r(F^{\prime})=r(F\cap F^{\prime})+r(F\cup F^{\prime})$ whenever $F$ and $F^{\prime}$ are flats. Assume that NN is a modular matroid. Then [Oxl11]*Theorem 11.4.10 tells us that we can obtain an amalgam of $M_{1}$ and $M_{2}$ by setting the rank of any subset $X\subseteq E_{1}\cup E_{2}$ to be
(1) $\min\{r_{1}(Y\cap E_{1})+r_{2}(Y\cap E_{2})-r_{N}(Y\cap\ell)\ \colon\ X\subseteq Y\subseteq E_{1}\cup E_{2}\}$
The resulting matroid is the proper amalgam of $M_{1}$ and $M_{2}$. We refer to the set $\ell$ as the amalgam base and we denote the proper amalgam by $\mathrm{Amal}_{\ell}(M_{1},M_{2})$, or $\mathrm{Amal}(M_{1},M_{2})$ if the amalgam base is clear from the context. We can easily check that every rank-22 matroid is modular. The following result is slightly more powerful than Proposition 4.1 in [MNW18] because it allows the matroids to be non-simple.
@@
We use NN to denote $M_{1}|\ell=M_{2}|\ell$. Assume $X\cap E_{1}$ is dependent in $M_{1}$. Then $X\cap E_{1}$ is dependent in $\mathrm{Amal}(M_{1},M_{2})$ since $\mathrm{Amal}(M_{1},M_{2})|E_{1}=M_{1}$. Thus $X$ is dependent in $\mathrm{Amal}(M_{1},M_{2})$. By symmetry we conclude that if $X\cap E_{1}$ is dependent in $M_{1}$ or if $X\cap E_{2}$ is dependent in $M_{2}$, then $X$ is dependent in $\mathrm{Amal}(M_{1},M_{2})$. Henceforth we assume $X\cap E_{1}$ is independent in $M_{1}$ and $X\cap E_{2}$ is independent in $M_{2}$.
@@
Let $Y$ be $X\cup y$. Because $y$ is in $\operatorname{cl}_{1}(X-E_{2})$ it follows that $r_{1}(X\cap E_{1})=r_{1}(Y\cap E_{1})$. Similarly, $r_{2}(X\cap E_{2})=r_{2}(Y\cap E_{2})$. If $X\cap\ell=\emptyset$ then $0=|X\cap\ell|=r_{N}(Y\cap\ell)-1$. Now assume that $X\cap\ell=\{x\}$. Since $\{x,y\}$ is independent in NN, we have $r_{N}(Y\cap\ell)=2$, so $|X\cap\ell|=r_{N}(Y\cap\ell)-1$ holds in either case. Now we see that
@@
Assume $y$ and $y^{\prime}$ are distinct elements of $Y\cap\ell$. If $\{y,y^{\prime}\}$ is dependent then we can assume without loss of generality that $y$ is not in $X$, or else $X\cap\ell$ is dependent in NN. Since $Y$ contains no loops, it follows that $\{y,y^{\prime}\}$ is a circuit of NN. This means that $y$ is in $\operatorname{cl}_{N}((Y-y)\cap\ell)$, and we have contradiction to Claim 4.1.1. Therefore $\{y,y^{\prime}\}$ is independent.
We know that $Y-X\subseteq\ell$ is non-empty so $Y\cap\ell$ contains at least one element. Let $y$ be such an element. Assume $Y\cap\ell$ contains three distinct elements and let $y_{1}$ and $y_{2}$ be distinct elements in $(Y\cap\ell)-y$. Then $\{y_{1},y_{2}\}$ is independent by the previous paragraph, so $\{y_{1},y_{2}\}$ spans NN. This means that $y$ is in $\operatorname{cl}_{N}((Y-y)\cap\ell)$, a contradiction. We conclude that $|Y\cap\ell|<3$. ∎
Assume $X\cap\ell$ contains distinct elements $x$ and $x^{\prime}$. Then $\{x,x^{\prime}\}$ is independent by Claim 4.1.2. In this case $\{x,x^{\prime}\}$ spans NN. We can let $y$ be an element in YXY-X, and now $y$ is in $\operatorname{cl}_{N}((Y-y)\cap\ell)$, contradicting Claim 4.1.1. Hence $|X\cap\ell|$ is in $\{0,1\}$.
@@
Assume that $X\cap\ell$ is non-empty, and therefore contains a single element $x$. Without loss of generality, we assume that $x=y^{\prime}$. Now $X\cap E_{1}=(Y-y)\cap E_{1}$, so $\operatorname{cl}_{1}(X\cap E_{1})$ contains both $x$ and $y$ by Claim 4.1.1. As $\{x,y\}$ is independent in NN by Claim 4.1.2, it follows that $\operatorname{cl}_{1}(X\cap E_{1})$ contains $\ell$, contradicting our earlier assumption. Therefore $X\cap\ell$ is empty and $Y\cap E_{1}=(X\cap E_{1})\cup\{y,y^{\prime}\}$.
@@
For $i=1,2$, let $\Omega_{i}=(G_{i},\mathcal{B}_{i})$ be a biased graph without balanced loops and assume that $V(G_{1})\cap V(G_{2})=\{u,v\}$. Assume also that in $\Omega_{1}$ and $\Omega_{2}$, every vertex is incident with an unbalanced loop, and that each vertex $w$ has two distinct neighbours $w_{1}$ and $w_{2}$ such that for each $i=1,2$, there is an unbalanced 22-edge cycle containing $w$ and $w_{i}$. Let $\Omega=(G,\mathcal{B})$ be a biased graph such that $G$ has no isolated vertices and $\mathrm{Amal}(F(\Omega_{1}),F(\Omega_{2}))$ is equal to $F(\Omega)$. Then $G$ is isomorphic to $G_{1}\oplus_{uv}G_{2}$.
@@
Let $L$ be a set containing exactly one unbalanced loop incident with each vertex in $G_{1}\oplus_{uv}G_{2}$. Then $L\cap E_{i}$ is a basis of $F(\Omega_{i})$ for each $i=1,2$. It is easy to verify, using Proposition 4.1, that $L$ is independent in $\mathrm{Amal}(F(\Omega_{1}),F(\Omega_{2}))$. Let $e$ be an arbitrary element of $\mathrm{Amal}(F(\Omega_{1}),F(\Omega_{2}))$ that is not in $L$. We can assume without loss of generality that $e$ is an edge of $\Omega_{1}$. Either $e$ is an unbalanced loop that has a common vertex with a loop in $L$, or it is non-loop edge of $\Omega_{1}$. In these cases, $e$ is in a 22- or 33-element circuit of $F(\Omega_{1})$ that is contained in $L\cup e$. This circuit is also a circuit of $\mathrm{Amal}(F(\Omega_{1}),F(\Omega_{2}))$. It now follows that $L$ spans $\mathrm{Amal}(F(\Omega_{1}),F(\Omega_{2}))$, and is therefore a basis of this matroid. Let $x$ and $y$ be an arbitrary pair of distinct vertices contained in an unbalanced 22-edge cycle of either $\Omega_{1}$ or $\Omega_{2}$. The edges from the 22-edge cycle along with unbalanced loops incident with $x$ and $y$ form a long line of either $F(\Omega_{1})$ or $F(\Omega_{2})$, and it is not difficult to see that this line is also a line in $\mathrm{Amal}(F(\Omega_{1}),F(\Omega_{2}))$. Now the hypotheses imply that each element of $L$ is in two distinct long lines of $\mathrm{Amal}(F(\Omega_{1}),F(\Omega_{2}))$. Now Proposition 2.1 implies that the elements of $L$ are all unbalanced loops in $\Omega$. Since $L$ is independent in $F(\Omega)$, we have no more than one element of $L$ incident with any given vertex of $G$, for any two loops incident with the same vertex form a handcuff. If there exists some vertex of $G$ not incident with a loop in $L$, then this vertex is incident with an edge $e$ (since $G$ has no isolated vertices), and now $e$ is not spanned by $L$. This is a contradiction, so every vertex of $G$ is incident with exactly one loop in $L$. This induces a bijection between the vertices of $G_{1}\oplus_{uv}G_{2}$ and $G$. Any edge of $G_{1}\oplus_{uv}G_{2}$ that is not in $L$ is incident with at most two elements of $L$, and this shows that the bijection between the vertices of $G_{1}\oplus_{uv}G_{2}$ and $G$ gives us an isomorphism between the two graphs. ∎
@@ 5. Uniformly locally finite groups
Figure 2. The graph $H_{n,N}\setminus H_{n,N}^{*}$.
@@
all in $\Gamma$ satisfying condition $(\dagger)$ in Definition 5.2. Let $\sigma_{N}$ be $\sigma(\Gamma,\{a_{1},\ldots,a_{n}\},s_{N},\mathfrak{M}_{N},\{d_{i,j}\},\{q_{i}\})$, the $\Gamma$-gaining of $H_{n,N}$. Then, for ease of notation, for each $N\geq 1$, let $H_{N}=(H_{n,N},\sigma_{N})$, and let $\Omega_{N}$ be the associated biased graph. Note that any $H_{N}$ contains $G_{\ell,N}$, and the $\{G_{\ell,N}\}$ are all isomorphic as $\Gamma$-gained graphs. So, we could arrange that the $H_{n}$ are all sub-graphs of some larger graph $\mathbb{G}$, and that the $G_{\ell,N}$ are actually equal, and for any $N,N^{\prime}$ distinct, $H_{N}\cap H_{N^{\prime}}=G_{\ell,N}$. Thus, given any two $H_{N},H_{N^{\prime}}$, we can construct both the matroid amalgam $\mathrm{Amal}(F(H_{N}),F(H_{N^{\prime}}))$ over the base $F(\ell_{N})=F(\ell_{N^{\prime}})$, and the graph amalgam $H_{n,N}\oplus_{\delta_{1},\delta_{2}}H_{n,N^{\prime}}$ over the base $\{\delta_{1},\delta_{2}\}=\{\delta_{1}^{\prime},\delta_{2}^{\prime}\}$. Throughout the proof, when we amalgamate it will always be over these bases. By Lemma 5.5, if the class of $\Gamma$-gain-graphic matroids is $\operatorname{CMSO}_{1}$-definable, there is a finite partition of $\omega$ such that for any $N,N^{\prime}$, whether $\mathrm{Amal}(F(H_{N}),F(H_{N^{\prime}}))$ is a $\Gamma$-gain-graphic matroid depends only on which classes in the partition NN and $N^{\prime}$ belong to. Thus, for a contradiction it suffices to show that the amalgam $\mathrm{Amal}(F(H_{N}),F(H_{N^{\prime}}))$ is a $\Gamma$-gain matroid if and only if $N=N^{\prime}$ for all $N,N^{\prime}\geq 1$. There are two directions to this.
@@
and conclude $\tau(B_{i,j})=\tau(A_{j})=h_{j}$. Now, let $\eta:\{1,\ldots,N\}\rightarrow\{h_{1},\ldots,h_{n}\}\cup\{\mathrm{Id}\}$ be a string of length at most NN in the $h_{i}$. There is a path $P_{\eta}$ from $\gamma_{1}$ to $\gamma_{N+1}$ such that for each $k$, the edge $P_{\eta}$ takes between $\gamma_{k}$ and $\gamma_{k+1}$ is mapped to $\eta(k)$ by $\tau$. Let $C_{\eta}$ be the cycle obtained from $P_{\eta}$ by appending $\gamma_{N+1}\beta_{2}\alpha_{2}\ A_{s}^{-1}\ \alpha_{1}\ K_{1}^{-1}\ \beta_{1}\ K_{2}\ \gamma_{1}$. Thus, $\tau(C_{\eta})=\tau(P_{\eta})h_{s}^{-1}$. Since $f_{\{a_{i}\}}(s)=N$, there exists $\eta$ a string of length NN such that $\sigma_{N}(C_{\eta})=\mathrm{Id}$, but not one of smaller length. So there exists $\eta$ of length NN such that $\tau(C_{\eta})=\mathrm{Id}$ but not one of smaller length. In other words, $f_{\{h_{i}\}}(h_{s})=N$. Symmetrically, if we let $\tau^{\prime}(C_{j})=h_{j}^{\prime}$ for all $j$, and $\tau^{\prime}(C_{s})=h_{s}^{\prime}$, then $f_{\{h_{i}^{\prime}\}}(h_{s}^{\prime})=N^{\prime}$. But the edges in $\mathcal{C}$ are all in the amalgamation base, so $\tau$ and $\tau^{\prime}$ agree on $\mathcal{C}$, so
@@
Now that we have this identity, we want to show $F(H_{n,N}\oplus_{\ell_{N}}H_{n,N^{\prime}},\sigma_{N}\cup\tau^{\prime})$ is the matroid $\mathrm{Amal}(F(H_{N}),F(H_{N^{\prime}}))=\mathrm{Amal}(F(H_{N}),F(H_{n,N^{\prime}},\tau^{\prime}))$, by Lemma $4.$ We must prove that this data satisfies its conditions. Note (i), (ii) are immediate from the definition of $\tau^{\prime}$, and (iii) is satisfied by $Q_{N+6}$ and $Q_{N+7}$, so there is only one condition of Lemma 4.4 remaining: to show that for any paths $P\subseteq H_{n,N}$, $P^{\prime}\subseteq H_{n,N^{\prime}}$, both starting at $\delta_{1}$ and ending at $\delta_{2}$, and $\sigma_{N}(P)=\tau^{\prime}(P^{\prime})$, there is some edge $E$ from $\delta_{1}$ to $\delta_{2}$ such that $\sigma_{N}(E)=\sigma_{N}(P)=\tau^{\prime}(P^{\prime})$. Fix such $P,P^{\prime}$. Their union is a balanced cycle, so they must lie in $H^{*}_{n,N}$ and $H^{*}_{n,N^{\prime}}$, respectively. Then, in particular, $f_{\{a_{i}\}}(P)=f_{\{a_{i}\}}(P^{\prime})$. By inspection $f_{\{a_{i}\}}(P)\leq 3N+1$. If $P^{\prime}$ passes through $K_{1}^{\prime}$ or $K_{2}^{\prime}$, $f_{\{a_{i}\}}(P^{\prime})\geq 3N+2$. Therefore, $P^{\prime}$ passes through neither $K_{1}^{\prime}$ nor $K_{2}^{\prime}$. Hence, $\tau^{\prime}(P^{\prime})\in\{\mathrm{Id},a_{1},\ldots,a_{n},s\}$, and so there is some edge $E\in\mathcal{A}$ such that $\sigma_{N}(E)=\sigma_{N}(P)=\tau^{\prime}(P^{\prime})$. The conditions of the Lemma are satisfied, so $\mathrm{Amal}(F(H_{N}),F(H_{N^{\prime}}))=F(H_{n,N}\oplus_{\ell_{N}}H_{n,N^{\prime}},\sigma_{N}\cup\tau^{\prime})$ is a $\Gamma$-gain graphic matroid. ∎